OpenStudy (b2st):

How to find a.... a^7 =7777. The problem is if a^7 = 7777 and (a^6)/(b) = 11 what is the value of ab.. The answer is apparently 707?? I don't understand how to get it. Please do it step by step please. Thank you! :D I'm stuck. Can

OpenStudy (psymon):

\[\frac{ a ^{6} }{ b }=11->a ^{6}=11b\] Multiply both sides by a \[a ^{7}=11ab\] But a^7 = 7777, so: \[7777=11ab\]

OpenStudy (b2st):

wow it was that easy. omg i'm really think way too hard for these problems... - o -

OpenStudy (b2st):

And thank you again! lol im going to fail the sat. >:

OpenStudy (psymon):

Lol, good luck. And np.

OpenStudy (wolf1728):

First, let's try to solve for 'a' a^7 = 7,777 7*log(a) = log(7,777) log(a) = 3.890812099/7 log(a) = 0.5558302999 10^0.5558302999 = a a = 3.5960879092 and a^7 = 7,777

OpenStudy (wolf1728):

wow - I was typing all that (and didn't get the final answer) as Psymon was typing out a much more concise answer Well I hope B2ST can use the value of 'a'

OpenStudy (b2st):

lol its okay.. I did similar to that.. - o - sat questions is too tricky...

OpenStudy (wolf1728):

glad you came back B2ST you pursued the less efficient route of taking logarithms of everything? LOL By the way I gave Psymon a medal for his elegant answer.

OpenStudy (psymon):

Lol, to be honest, I wouldnt have even thought of brute-force solving for a. I figured there would just be a way to relate all of the given information. I solved for a^6 and then just saw what to do.

OpenStudy (wolf1728):

I think "brute force" and stick with it. (Not always a good thing to do). But heck, we all know the value of 'a' now.

OpenStudy (psymon):

Lol, nvm xD

OpenStudy (b2st):

I'm taking these questions way too complicated. ; A; I cant simplify my mind anymore. lol

OpenStudy (wolf1728):

if these are SAT practice questions they will purposely throw in problems that have a difficult solution AND an easy one. Example: Which is larger? a) 237/5,693 or b) 237/5,694 One solution takes the time the other solution is much quicker.

OpenStudy (b2st):

I'm using Baron's math work book. So the questions are more difficult than the SAT..

OpenStudy (wolf1728):

Well it's good to see you are not taking the easy route. LOL